OpenStudy (anonymous):

solve the system 2x+3y=6, x-3y=5 by graphing. And is there a solution?

OpenStudy (alexwee123):

try substituting

OpenStudy (precal):

do you have a graphing calculator?

OpenStudy (precal):

x=3.666666 y=-.444444

OpenStudy (alexwee123):

2x+3y=6, x-3y=5 x=3y+5 plug x into 2x+3y=6 to find out the y coordinate of the intersection

OpenStudy (precal):

x=(11/3) y=(-4/9)

OpenStudy (alexwee123):

then plug in the y value after solving out into either equation to find the x value of the intersection

OpenStudy (anonymous):

is it x=0 and y=2

OpenStudy (precal):

no it is x=11/3 and y=-4/9

OpenStudy (anonymous):

okay

OpenStudy (precal):

do you have a graphing calculator?

OpenStudy (anonymous):

no

OpenStudy (precal):

https://www.desmos.com/calculator here is a free online

OpenStudy (anonymous):

k thanks

OpenStudy (precal):

anytime :)